### v2

@@ 4.3 The Kirkwood-closure process
A point process ${\mathord{\mathbb{P}}}$ is called the Kirkwood-closure process for a given density $\rho$ and a non-negative even function $g$ on ${\mathord{\mathbb{R}}}^{d}$ if the correlation functions of ${\mathord{\mathbb{P}}}$ satisfy
@@
It was shown in [18] by Kuna et al. that, if $C(g)=\int_{{\mathord{\mathbb{R}}}^{d}}|g(x)|\,\mathrm{d}x<\infty$ and there is a $b\geq 1$ such that $\prod_{i=1}^{n}g(x_{i}-x_{0})\leq b$ for any $x_{0},x_{1},\dots,x_{n}\in{\mathord{\mathbb{R}}}^{d}$ with $\prod_{i<j}g(x_{i}-x_{j})>0$, then the Kirkwood-closure process exists for $\rho<(ebC(g))^{-1}$. This was an extension of the results in [1], where only the case $g\leq 1$ was discussed. The Kirkwood-closure, which in computational physics is often called Kirkwood-superposition was first suggested in [16] and has been widely used to approximate higher-order correlation functions and thus just having to calculate $\rho$ and $\rho^{(2)}$ from simulation data. In this case, it also known that ${\mathord{\mathbb{P}}}$ satisfies a Ruelle condition with $\xi=\rho b^{1/2}$ and for every bounded $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$
$\sup_{x\in{\mathord{\mathbb{R}}}^{d}}\int_{\Lambda^{n}}\left|{\rho}_{T}^{(1+n)}(x,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\leq n!D\left(\rho ebC(g)\right)^{n},$
@@
for $x\in{\mathord{\mathbb{R}}}^{d}$ and $\eta\in\Gamma$. Note that, due to the conditions on $g$ the function $k$ is well-defined on ${\mathord{\mathbb{R}}}^{d}\times\Gamma$ und mild decay conditions on $g$. Since $k(x,\eta)=\exp(\mu-W(\{x\}\mid\eta))$, i.e. $k(x,\emptyset)=e^{\mu}$, it thus follows that $\mu$ is given by
@@
Let $g$ satisfy the assumptions above and ${\mathord{\mathbb{P}}}$ be the Kirkwood-closure process if $\rho<((2+\zeta D)ebC(g))^{-1}$ then the chemical potential $\mu$ of the Kirkwood-closure process is given by (10).
@@
5 Motivation: The case of the zero-point density
@@
$\int_{\Lambda}\int_{\Lambda^{n}}\left|{\rho}_{T}^{(1+n)}(x,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\,\mathrm{d}x\leq|\Lambda|n!Dq^{n}.$
This means the assumptions of Theorem A are satisfied and we can write
@@
Let us consider the case that $H$ consists only of a pair interaction, i.e. a measurable even function $u\colon{\mathord{\mathbb{R}}}^{d}\to{\mathord{\mathbb{R}}}\cup\{+\infty\}$ such that $H(\{\textit{{x}}_{n}\})=\sum_{i<j}u(x_{i}-x_{j})$. If we further assume that $u$ is superstable, i.e. that there exists $r_{0}>0$ and decreasing positive functions $\varphi:(0,r_{0})\to{\mathord{\mathbb{R}}}^{+}_{0}$ and $\psi:[0,\infty)\to{\mathord{\mathbb{R}}}^{+}$ with
and
$u(x)$ $\,\geq\,\varphi(|x|)\qquad\textrm{ for }\ 0<|x|<r_{0}\,,$
$|u(x)|$ $\,\leq\,\psi(|x|)\qquad\textrm{ for }\ |x|\geq r_{0}\,,$
then (8) is satisfied if
$e^{\mu}=\bar{q}\left(e^{2B+1}C(u)\right)^{-1}\qquad\text{ for some }\qquad\bar{q}<1/3,$ (16)
where $C(u):=\int_{{\mathord{\mathbb{R}}}^{d}}|e^{-u(x)}-1|\,\mathrm{d}x$ and $B=B(u)>0$ is the stability constant of $H$, with
$D=\frac{e^{\mu}e^{-2B}}{(1-\bar{q})}\qquad\text{ and }\qquad q=\frac{\bar{q}}{1-\bar{q}},$
cf. [25]. In this case in particular taking the thermodynamic limit of (15), i.e. dividing by the volume $|\Lambda|$ and increasing $\Lambda$ to the whole space ${\mathord{\mathbb{R}}}^{d}$, gives
@@
where $p$ is the infinite volume grand canonical pressure, cf. [23]. The above expression follows immediately from the cluster expansions of the pressure and the truncated correlation functions therein and [24].
@@ 6 Generalization to the one-point density
We now want to generalize this approach for the one-particle density $j^{(1)}_{\Lambda}$ of an infinite volume Gibbs measure. Again we will use (2) to find an expansion and (1) to identify the terms. By (2) we have
@@
Finally we plug the above into (21) to get
@@
Finally using (11) proves (19).
@@
$\int_{\Lambda^{k}}\widetilde{\rho}_{T}^{\,(1+k)}(x,\textit{{y}}_{k})\,\mathrm{d}\textit{{y}}_{k}=\int_{\Lambda^{k}}\frac{\rho^{(1+k)}_{T}(x,\textit{{y}}_{k})}{\rho}\,\mathrm{d}\textit{{y}}_{k}-\sum_{l=2}^{k}\sum_{\pi\in\Pi_{l}(k)}\prod_{i=1}^{l}\int_{\Lambda^{\kappa_{i}}}\widetilde{\rho}^{\,(1+\kappa_{i})}_{T}(x,\pi_{i})\mathop{}\!\mathrm{d}{\pi_{i}}.$
Note that because we integrate over all $\textit{{y}}_{k}$, when looking at a particular partition $\pi\in\Pi_{l}(k)$ it does not matter which particular elements of $\textit{{y}}_{k}$ are contained in which part $\pi_{i}$ of $\pi$ only the sizes of the different parts $\pi_{i}$ matter. This is the definition of the $k$th Bell polynomial, cf. [7], and allows us to write
@@
The claim is thus proved and we arrive at the bound
@@
Proof of Theorem 3.1. Assume that ${\mathord{\mathbb{P}}}$ is a $(\mu,H)$-Gibbs measure satisfying assumptions A and B. By (19), the triangle inequality and Theorem 6.3 we get for $F_{n}$ of (18), that
$\int_{\Lambda^{n}}\left|F_{n}(0,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\leq|\Lambda|(n-1)!Dq^{n-1}+M(n-1)!(1+\zeta D)^{n}q^{n}\leq\left(|\Lambda|\frac{D}{q}+M\right)(n-1)!\left(q(2+\zeta D)\right)^{n}.$
@@
By assumption A the left-hand side of (33) converges to $\mu$ as $\Lambda\nearrow{\mathord{\mathbb{R}}}^{d}$ and since the bound in (23) is independent of $\Lambda$ the right-hand side above converges by dominated convergence and we arrive at (10).

### v3

@@ 4.3 The Kirkwood-closure process
A point process ${\mathord{\mathbb{P}}}$ is called the Kirkwood-closure process for a given density $\rho>0$ and a non-negative even function $g$ on ${\mathord{\mathbb{R}}}^{d}$ if the correlation functions of ${\mathord{\mathbb{P}}}$ satisfy
@@
It was shown in [18] by Kuna et al. that, if $C(g):=\int_{{\mathord{\mathbb{R}}}^{d}}|g(x)-1|\,\mathrm{d}x<\infty$ and there is a $b\geq 1$ such that $\prod_{i=1}^{n}g(x_{i}-x_{0})\leq b$ for any $x_{0},x_{1},\dots,x_{n}\in{\mathord{\mathbb{R}}}^{d}$ with $\prod_{i<j}g(x_{i}-x_{j})>0$, then the Kirkwood-closure process exists for $\rho<(ebC(g))^{-1}$. This was an extension of the results in [1], where only the case $g\leq 1$ was discussed. The Kirkwood-closure, which in computational physics is often called Kirkwood-superposition was first suggested in [16] and has been widely used to approximate higher-order correlation functions and thus just having to calculate $\rho$ and $\rho^{(2)}$ from simulation data. In this case, it also known that ${\mathord{\mathbb{P}}}$ satisfies a Ruelle condition with $\xi=\rho b^{1/2}$ and for every bounded $\Lambda\subset{\mathord{\mathbb{R}}}^{d}$
$\sup_{x\in{\mathord{\mathbb{R}}}^{d}}\int_{\Lambda^{n}}\left|{\rho}_{T}^{(1+n)}(x,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\leq n!D\left(\rho ebC(g)\right)^{n+1},$
@@
for $x\in{\mathord{\mathbb{R}}}^{d}$ and $\eta\in\Gamma$. Note that, due to the conditions on $g$ the function $k$ is well-defined on ${\mathord{\mathbb{R}}}^{d}\times\Gamma$ and mild decay conditions on $g$. Since $k(x,\eta)=\exp(\mu-W(\{x\}\mid\eta))$, i.e. $k(x,\emptyset)=e^{\mu}$, it thus follows that $\mu$ is given by
@@
Let $g$ satisfy the assumptions above and ${\mathord{\mathbb{P}}}$ be the Kirkwood-closure process if
$\rho<\frac{1}{(2+\zeta D)ebC(g)}$
then the chemical potential $\mu$ of the Kirkwood-closure process is given by (9).
Remark 4.7.
In the case of the Kirkwood-closure it is quite easy to see the connection of the expansion (9) to classic cluster expansions. It follows from (14) and Section 3.1 [18] that
$\widetilde{\rho}_{T}^{\,(1+k)}(\tilde{\textit{{y}}}_{k+1})=\rho^{k}\sum_{C\in\mathcal{C}_{k+1}}\prod_{\{i,j\}\in E(C)}\left(g(\tilde{y}_{i}-\tilde{y}_{j})-1\right)$
where $\tilde{\textit{{y}}}_{k+1}=(x,\textit{{y}}_{k})$. Thus, in this case we have
$\mu=\log\rho+\sum_{k=1}^{\infty}\frac{(-\rho)^{k}}{k!}\int_{({\mathord{\mathbb{R}}}^{d})^{k}}\sum_{C\in\mathcal{C}_{k+1}}\prod_{\{i,j\}\in E(C)}\left(g(\tilde{y}_{i}-\tilde{y}_{j})-1\right)\,\mathrm{d}\textit{{y}}_{k},$
where $\mathcal{C}_{k+1}$ is the set of connected graphs on $k+1$ vertices and $E(C)$ is the set of edges of the graph $C$. This has the same structure as the classic cluster expansion for the density in the case of pair-interactions with $-\rho$ taking the role of the activity (the exponential of the chemical potential) and $g-1$ taking the role of the Mayer function.
@@
5 A first step: The case of the zero-point density
@@
$\int_{\Lambda}\int_{\Lambda^{n}}\left|{\rho}_{T}^{(1+n)}(x,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\,\mathrm{d}x\leq|\Lambda|n!D\rho q^{n}.$ (15)
This means the assumptions of Theorem A are satisfied, and we can write
@@
Let us consider the same setting as in Example 4.1. If in this case in (11) we have $\bar{q}<1/3$, then $q=\bar{q}/(1-\bar{q})<1/2$ and thus taking the thermodynamic limit of (16), i.e. dividing by the volume $|\Lambda|$ and increasing $\Lambda$ to the whole space ${\mathord{\mathbb{R}}}^{d}$, gives
@@
where $p$ is the infinite-volume grand-canonical pressure, cf. [23]. The above expression follows immediately from the cluster expansions of the pressure and the truncated correlation functions therein and [24].
@@ 6 Generalization to the one-point density
We now want to generalize this approach for the one-particle density $j^{(1)}_{\Lambda}$ of an infinite-volume Gibbs measure. Again we will use (2) to find an expansion and (1) to identify the terms. By (2) we have
@@
Finally, we plug the above into (21) to get
@@
Finally, using (10) proves (19).
@@
$\int_{\Lambda^{k}}\widetilde{\rho}_{T}^{\,(1+k)}(x,\textit{{y}}_{k})\,\mathrm{d}\textit{{y}}_{k}=\int_{\Lambda^{k}}\frac{\rho^{(1+k)}_{T}(x,\textit{{y}}_{k})}{\rho}\,\mathrm{d}\textit{{y}}_{k}-\sum_{l=2}^{k}\sum_{\pi\in\Pi_{l}(\{\textit{{y}}_{k}\})}\prod_{i=1}^{l}\int_{\Lambda^{\kappa_{i}}}\widetilde{\rho}^{\,(1+\kappa_{i})}_{T}(x,\pi_{i})\mathop{}\!\mathrm{d}{\pi_{i}}.$
Note that because we integrate over all $\textit{{y}}_{k}$, when looking at a particular partition $\pi\in\Pi_{l}(\{\textit{{y}}_{k}\})$ it does not matter which particular elements of $\textit{{y}}_{k}$ are contained in which part $\pi_{i}$ of $\pi$ only the sizes of the different parts $\pi_{i}$ matter. This is the definition of the $k$th Bell polynomial, cf. [7], and allows us to write
@@
The claim is thus proved, and we arrive at the bound
@@
Proof of Theorem 3.1. Assume that ${\mathord{\mathbb{P}}}$ is a $(\mu,H)$-Gibbs measure satisfying a Ruelle-condition as well as Assumptions A and B. By (19), the triangle inequality, (15) and Theorem 6.3 we get for $F_{n}$ of (18), that
$\int_{\Lambda^{n}}\left|F_{n}(0,\textit{{y}}_{n})\right|\,\mathrm{d}\textit{{y}}_{n}\leq|\Lambda|(n-1)!D\rho q^{n-1}+M(n-1)!(1+\zeta D)^{n}q^{n}\leq\left(|\Lambda|\frac{D\rho}{q}+M\right)(n-1)!\left(q(2+\zeta D)\right)^{n}.$
@@
By Assumption A and (1) the left-hand side of (33) converges to $\mu$ as $\Lambda\nearrow{\mathord{\mathbb{R}}}^{d}$ and since the bound in (23) is independent of $\Lambda$ the right-hand side above converges by dominated convergence, and we arrive at (9).